# Optimizing energy conversion with nonthermal resources in steady-state quantum devices

### v1

@@ Appendix D Numerical procedure
Several results presented in Sec. IV (Figs. 7 and 8 in particular) are calculated fully numerically. This Appendix summarizes the numerical procedure that was used to obtain these results. As per the data availability statement, the code and data used are all freely available and further details can be found there [62].
The code for producing the data is written in Python and the numerical method relies heavily on the SciPy module [71] for equation solvers and integrations. Principally, the challenge of applying the methods in Sec. III lies in the Heaviside functions; because large parts of the energy spectrum may be zero, especially for low fixed currents, standard integration methods can produce erroneous results, usually returning zero. Instead, the integration is divided into several areas, defined by the roots of the condition inside the Heaviside functions. A standard integration is made over each of these areas and added together to form the calculated current or noise. This ensures accurate results for any optimal transmission function, especially correcting for cases like Fig. 7(n), where the boxcars are narrow and disconnected. The roots calculation does add some computational complexity, but the impact is small in our implementation despite the brute-force minimization over the energy spectrum.

### v2

@@ Appendix D Numerical procedure
Several results presented in Sec. IV (Figs. 7 and 8 in particular) are calculated fully numerically. This Appendix summarizes the numerical procedure that was used to obtain these results. As per the data availability statement, the code and data used are all freely available and further details can be found there [zenodo].
The code for producing the data is written in Python and the numerical method relies heavily on the SciPy module [2020SciPy-NMeth] for equation solvers and integrations. Principally, the challenge of applying the methods in Sec. III lies in the Heaviside functions; because large parts of the energy spectrum may be zero, especially for low fixed currents, standard integration methods can produce erroneous results, usually returning zero. Instead, the integration is divided into several areas, defined by the roots of the condition inside the Heaviside functions. A standard integration is made over each of these areas and added together to form the calculated current or noise. This ensures accurate results for any optimal transmission function, especially correcting for cases like Fig. 7(n), where the boxcars are narrow and disconnected. The roots calculation does add some computational complexity, but the impact is small in our implementation despite the brute-force minimization over the energy spectrum.